### v1

# Artificial Human Intelligence: The role of Humans in the Development of Next Generation AI

Abstract
Human intelligence, the most evident and accessible form of source of reasoning, hosted by biological hardware, has evolved and been refined over thousands of years, positioning itself today to create new artificial forms and preparing to self–design their evolutionary path forward. Beginning with the advent of foundation models, the rate at which human and artificial intelligence interact with each other has surpassed any anticipated quantitative figures. The close engagement led to both bits of intelligence to be impacted in various ways, which naturally resulted in complex confluences that warrant close scrutiny. In the sequel, we shall explore the interplay between human and machine intelligence, focusing on the crucial role humans play in developing ethical, responsible, and robust intelligent systems. We slightly delve into interesting aspects of implementation inspired by the mechanisms underlying neuroscience and human cognition. Additionally, we propose future perspectives, capitalizing on the advantages of symbiotic designs to suggest a human-centered direction for next-generation AI development. We finalize this evolving document with a few thoughts and open questions yet to be addressed by the broader community.
Keywords: Human-inspired AI, neuroscience, human vision, computer vision, cognition, neural networks, biological plausibility, CNN, KAN, Transformers.
1 Introduction
In the wake of the artificial intelligence (AI) revolution, industries and societies, more broadly almost everyday life errands have undergone profound transformations. The relentless evolution of AI technologies and digital learning mechanisms prompts critical inquiries into the extent of human contribution in steering these advancements. While modern AI systems showcase unparalleled prowess in automation, decision-making, and closed-domain problem-solving, the indispensability of human engagement and intervention remains unequivocal in shaping the trajectory of next-generation AI systems as it has been the source of inspiration since the very early appearances in the past [1].
The complexity of the intelligence problem, whose definition is yet to be established, is tempting to adapt human model in search of any type of intelligent behavior. Paradoxically, any conceivable form of artificial intelligence will inevitably originate either by mimicking human form or be a product of human intellect, at least in the near term, unless artificially generated forms begin building up newer forms of intelligences. Consequently, bias might be an intrinsic part of its initial design. However, implementation constraints, task variability, and interactions with other intelligent forms could guide humans in creating general intelligence to collaborate with [2]. Moreover, this collaboration could guide humanity to discover new inventions or new cognitive technologies, which might potentially expand the range of human thought. This virtuous feedback could eventually lead to artificially generated human-like intelligence, a byproduct of symbiotic evolution of digital and biological computational models.
Figure 1: Summary of the manuscript detailing various topics of interest and where they are covered with associated section and/or subsection numbers.
In this brief exposition, our objective is to investigate the strategic placement of human–centered inquiries in the advancement of contemporary AI systems despite the recent remarkable (not necessarily human–centric) developments in machine learning and computer vision fields [3]. Considering the strategies of machine and human intelligences in completing various tasks and their complementarity in behaviour, a potential fusion of both emerge as a plausible option that might soon become an applicable reality. Since human-centricity seems to be unavoidable in any type of AI-related research, we have decided to label this specific fusion of intelligences and all the interactions within as “artificial human intelligence” and explicate various intriguing topics of discourse for the mutually beneficial association of the two (seemingly) mechanistically distinct notions. Naturally, the exploration of both forms of intelligence necessitates addressing numerous intriguing questions. These include early human brain developments and their connections to neuroscience, as well as biological hardware dependencies, efficiency of implementations and the potential long-term effects on both human and machine intelligences as both intelligent forms will likely coexist in any foreseeable future.
A summary of the paper’s organization is briefly illustrated in Fig. 1. The paper begins by exploring the necessity for different definitions of intelligence, provide some of the expert opinions on the subject and its origins which later is followed by the introduction of the novel notion of artificial human intelligence, arguing in favor of the inevitable confluence of artificial and natural intelligences. This unique interaction, we hypothesize, would be the defacto standard of future human-to-human as well as human-to-machine communications and interfaces. It must be emphasized though that such an interaction is not an alternative to Artificial General Intelligence (AGI) [4, 5, 6] in any sense, but rather a tentative direction achieving more generic and capable machine intelligence with arguably so many human-like features at behavioral, architectural and functional levels [2]. Later, we provide details about what we mean by artificial human intelligence and review few attempts as the state-of-the-art (SoTA) in this field. We further use brain-centric categorization of past efforts and consider the missing elements in brain-inspired information processing. We delve into critical questions that concern where exactly humans will be positioned in the newly formed artificial intelligence ecosystem, addressing few challenges and discuss the final alignment with humans. Finally, we comment/speculate on the effects of AI on the human brain and cognition and the set of required management actions in the long term as part of our discussions.
2 Views on Intelligence
The notion of intelligence, although seemingly straightforward in everyday language to define or use, is remarkably sophisticated and complex. Attributes like expeditious learning or life-long adaptation, knowledge accumulation and representation, abstract reasoning, symbolic problem-solving, and creativity contribute to its elusive and unfathomable nature. Psychologists have long debated its definition and measurement, sparking controversies regarding human valuation and performance assessment, particularly relevant and applicable to current education system [7]. Despite some consensus within the scientific community on measuring intelligence via standarized tests [8], issues persist around their scope, potential biases, and the spectrum of skills they believe these experiments assess. Concurrently, the discourse on machine intelligence faces even greater complexity due to the diverse forms and abilities of machines. Defining intelligence in precise terms, applicable across a wide range of systems, poses significant challenges, especially as tasks once considered benchmarks of human intellect become now remarkably trivial for machines (like ImageNet [9] or EcoSet [10] based categorization tasks). As technology evolves, so too does our perception of intelligence, necessitating a comprehensive and enduring definition that transcends sensory limitations, environmental constraints, and specific hardware types, yet remains practically achievable.
The quest for defining intelligence is a multifaceted challenge, labyrinthine in its nature and diverse in its interpretations. Intelligence includes a spectrum of cognitive abilities, adaptive behaviors, and problem-solving skills exhibited predominantly by living organisms. At its center, intelligence refers to the capacity to acquire knowledge, reason abstractly (including counterfactual thinking), solve problems, learn from exploratory experience (through interactions with the three dimensional world model), and adapt to unseen circumstances through generalization, quickly without necessarily using many examples (one-shot learning [11]). Intelligence might manifest itself in diverse forms across species, each tailored to suit the ecological niche and evolutionary context of the organism. Human intelligence, often pointed at as the cornerstone of cognitive evolution, is characterized by its complexity, versatility, and capacity for symbolic thought, language, and cultural and societal intercourse. However, defining human intelligence as the ultimate form of intelligence risks anthropocentric bias and overlooks the diverse manifestations of intelligence found in other organisms [12] or more recently in self-learning machines [13].
Scholars have proposed various models and frameworks to elucidate the multidimensional nature of intelligence. Gardner’s theory of multiple intelligences [14] posits that the broad notion of intelligence should comprise distinct modalities, including linguistic, logical-mathematical, spatial, musical, bodily-kinesthetic, inter-personal, intra-personal, and naturalistic intelligences. Each intelligence represents a domain of expertise and proficiency, offering insights into the diverse talents and capabilities of individuals. On the other hand, recent developments in deep learning have enabled rather constrained but various forms of intelligences capable of performing complex tasks and sometimes exhibiting human-like cognitive abilities [15]. This pluralistic account of intelligence allows a rethinking of developmental trajectory with necessary cross modality interactions. However, replicating the full breadth and depth of human intelligence remains elusive, as deep learning systems often excel in narrow domains but struggle with generalization, common-sense reasoning, and context-dependent understanding. Furthermore, ethical considerations surrounding the learning systems raise concerns about the societal impact of new forms of intelligences, including issues of autonomy, accountability, and bias. As we venture into the realms of various forms of artificial and/or non-human intelligence, navigating these challenges requires careful consideration of ethical, social, and philosophical implications.
On a lower scale, Brooks [16] highlights that in the early stages of AI research, intelligence was defined as tasks that highly educated scientists found challenging, such as chess, symbolic integration, proving mathematical theorems, and solving complex word algebra problems [17]. Tasks that young children effortlessly accomplished, like visually distinguishing between objects like a coffee cup and a chair, walking on two legs without losing balance, or navigating from their bedroom to the living room, were not initially regarded as activities requiring intelligence. Thus, he argues that a more informal notion of intelligence should be the sort of staff that humans do, pretty much all the time. Winston [18] delineates the objectives of Artificial Intelligence as twofold: the development of practical intelligent systems and the comprehension of human intelligence, arguing in favor of a mutually beneficial enterprise. On the other hand, Legg and Hutter [19] argues that despite the diverse set of definitions, it might still be possible to construct a formal encompassing definition of intelligence (e.g. universal intelligence), which is believed to have strong connections to the theory of optimal learning agents [20]. Optimal learning agents are artificial agents designed to learn and make decisions in a way that maximizes their performance or efficiency in achieving specific goals or tasks. These agents utilize various algorithms and strategies to improve their learning process and decision-making abilities over time through interaction and continuous feedback.
Studying the intelligence and its critical role in the evaluation of society is yet another challenge to face. Intelligence tests are typically used to provide summative scores for the assessment of academic and vocational success. However, it is argued that these tests could be limiting our understanding of intelligence. For example, the absence of precise and systematic scoring methods that leverage the multiple contributions of a response is one of the ways in which the current use of psychometric tests could be limiting our understanding of intelligence [21]. This limitation is believed to be due to the ways we design and execute tests and handle the test outcomes later for analysis.
3 Origins and the Path leading to AHI
Neither the origin of intelligence nor a theory for this origin have been adequately considered and addressed by the past literature. In fact, all definitions of intelligence seem to have a strong subjective component. Whether it be Turing or Newell [22], their descriptions of intelligence, involving knowledge-level considerations, beliefs, and intentions to a system, is strictly based on human judgements making both of them subjective. On the contrary, Steel [23] supports an alternative definition which highlights the continuum nature of it which progressively evolves from chemical systems to actual living organisms. Through precise definition of evolving complex adaptive systems, the quest for origins is reduced down to addressing two fundamental questions, namely, (1) how the brain demonstrates its remarkable behavioral plasticity and (2) how a symbolic layer might have emerged early in the development [23, 24, 25].
Based on our past human experience and historical account that intelligence most likely emerged from active perception and tight/attentive mechanisms that entail 3D interactions with the real-world phenomena through multi-sensory input [26, 27]. In other words, the biological intelligence requires to exist in multi-dimensional feedback-enabled open environments to emerge and develop [27]. We can think of different models of natural intelligence [14] [^2] According to Gardner, natural(istic) intelligence refers to the ability to recognize, categorize, and interact with elements of the environment, including objects, animals, and plants. This type of intelligence enables us to distinguish between different species, groups of people, or objects and to comprehend their interrelationships., but one of the observable models that is immediately accessible and intriguing is the human intelligence which has been leveraged by machine intelligence innovators (particularly deep learning community) in the past for so many years to inspire novel architectures, training methodologies, learning techniques, etc. Multi-model sensory input processing is essential to human intelligence as there is considerable evidence to support their joint progression early in childhood [28, 29]. In fact, recent work on the computational modellings of visually guided haptic sensory systems advocate for the value of multimodal fusion to improve the visual experience and performance [30]. As different sensory systems can be linked to distinct intelligent behaviors, growing number of studies seem to suggest that their synchronization and fusion lead to more accuracy in completion of specific tasks [31], which demonstrates the potential for integrating more general and non-biological intelligences. One of the most compelling pieces of evidence supporting this finding is the success of multimodal foundation models, which have demonstrated that a missing modality can be compensated for by the presence of other modalities as part of their emergent behavior [32].
Inspired from the past studies, we characterize AHI to describe trichotomous, consisting of human-assisted, human-inspired or human-independent machine intelligences. Since each one of these trejectories are rooted based on human capabilities and human intelligence, we coined the name AHI to describe all. Just like found in the integration of multimodal sensory inputs, this paradigm’s objective is to improve system performance by leveraging the distinct learning mechanisms of both humans and various AI approaches, which might result in complementarity in many aspects of processing and eventually in the final task performance. Additionally, the involvement of humans or understanding tasks performed by the human brain can help regulate ethical considerations, symbolic computations, and emotional aspects of the underlying functional systems.
Figure 2: In human-inspired intelligence, developments in machine intelligence are primarily driven by insights from biology and neuroscience, leading to overlapping advancements in both fields as each one of them evolves over time. Conversely, in human-assisted intelligence, the overlap tends to diminish over time, resulting in more specialized assistance from either humans or machines, thereby fostering the development of hybrid systems. Finally, in human-independent intelligence, the intrinsic features of machine intelligence evolve freely, naturally causing a divergence in the fundamental principles governing human and machine intelligence as the constraints such as hardware that hold up these intelligences become more diverse.
3.1 The current SoTA and Taxonomy
Current AI systems do not allocate enough emphasis to the definition of the intelligence problem, nor they seem to necessarily align in their objectives with different forms of human and/or non-human intelligences. Most early literature was inclined to adopt the model of human intelligence from a pragmatic perspective: The layered models of mind [33] (see also AlexNet [34] or VGGNet [35] etc.), a representation of layered cortical structures and hierarchical activations of neural populations, inspired inductive biases which eventually got encoded into the model architectures. They typically define the elevation to this intelligence as getting the job done faster, more efficient and accurate than most humans do. While engineering the machines, the evolution of the intelligence (and its definition) will not necessarily be following the same trajectory and hence be compatible with the model of human intelligence and its development. Although many of the most popular architectural models, such as residual networks [36, 37, 38, 39, 40], primarily designed to address a hardware or an algorithm issue, later has been shown to possess biological parallels both physiologically and functionally[15]. There are already many attempts in literature exploring this implicit comparison with respect to human systems [41, 42, 43].
In this work, we hypothesize that three branches of research can be leveraged to develop AI models in observation of human intelligence. One that could directly be inspired by the human brain, putting emphasis on its structural and functional connectivity, elements of developmental trajectory, and learning mechanisms. Secondly, different machine models can assist each other to create a more collective intelligence which will naturally lead to a human-assisted rather than a human-inspired model development. In this paradigm, models could be developed in such a way that functional complementarity is ensured based on the behavioral aspects of human intelligence as a reference. Finally, models can be developed independently, distributed emergence of independent intelligences, exploring freely the novel ways of task completions and discoveries of new implementation techniques that would otherwise be impossible to implement on biological hardware.
Figure 3: Categorization of Human-centered intelligences and associated subcatagories.
3.2 Human-inspired, Human-assisted, Human-Independent AI
The distinction between “human-inspired” and “human-assisted” AI pertains to the extent of human involvement and influence in the development and operation of artificial intelligence systems. A summary of this taxonomy is illustrated geometrically in Fig. 2, along with the temporal progression (directed arrows) of each paradigm in the evolutionary trajectory of the work that has been done in the last few decades (from Past to Future). More specifically, assuming stable centroids of the studies conducted around human and machine intelligences, a human-inspired trajectory will expand future works around these centroids, thereby naturally leading to more commonalities, i.e., increased biological plausibility due to everlasting incremental research around the same or similar ideas. In the human-assisted approach, the trend is reverse i.e., the future overlap between machine and human intelligences tend to shrink allowing hybrid designs to emerge and have better capacity to effectively address intelligence problem. Human assistance is not used as part of the training processs but rather used as the essential component of the system. On the contrary, human-independent trend moves the centroids of the literature works in different directions while their expansion is likely to continue, potentially creating more overlaps between the two distinct intelligences. Fig. 3 illustrates these options in detail and in the following subsections, we shall closely explore verbally each one of these options.
3.2.1 Human–inspired AI: Neuroscience meets Machines
In Human-inspired AI, machines are designed to mimic or replicate aspects of human intelligence, behavior, or complex cognitive processes. Part of this endevour could be inspired by the brain mechanisms responsible for processing information and navigating attention to complete complex tasks. Human-inspired AI aims at understanding and replicating the neural mechanisms behind human cognitive processes, such as learning, common sense reasoning, perception, and critical decision-making. This approach often involves studying biological systems closely including neuroscience, psychology, cellular chemistry and cognitive sciences to accurately inform and guide the design of AI algorithms and architectures. Perhaps the most critical phase of the formulation of human–like intelligence is the developmental phases of the brain anatomy and its training (learning trajectory) with psychophysical limitations impacting its evolution. Examples include neural networks inspired by the brain’s architecture (structural and functional connectivity) and specialization of the specific human brain regions, biologically plausible learning algorithms [44] or algorithms (recursive, top-down and bottom-up) that accurately model human decision-making processes. Human-inspired AI can be designed through methodological changes in model architecture or training. Architectural changes could be implemented through structural or functional changes involving addition or removal of nuronal units in the process of training/learning processes.
Figure 4: An example of biomimetic progreession of image resolution and color cues in the early months of development. Color maturization happens based on the wavelength of light i.e., red first accompanied by orange, followed by yellow, green, blue and purple towards the end of the spectrum.
Machines can be inspired by humans in various conspicuous ways. For instance, humans are recently shown to perform blur-detection tasks on faces using two different dominant strategies [45]. These strategies are under the influence of various factors including but not limited to top-down processes, bias, and behavioral similarities or differences. Behavioral characterizations of such human proclivities [46, 47] can be quantified and transferred to machine implementations to advance their guessing strength for multiple-choice tasks. Guessing framework is mathematically sound and inspiring piece of tool to design as well as evaluate neural network strategies with the potential of optimal guessing techniques being incorporated as part of learning [48, 49]. Moreover, psychophysics and neural activations can be used to understand the temporal ordering of critical processes and spatial distribution of basic functions, such as perception, performed by specialized cortical circuits. The results of said studies are essential to choosing the right inductive biases, architecture as well as learning (optimization) schemes for embedding human decision-making processes as part of the model development. It is quite a fortune that the human brain can region-specifically be monitored while actively functioning through information extraction techniques based on electrical or magnetic principles. The neuroimaging and electrophysiological methodologies that come in handy in the analysis of the human brain include EEG [50], EMG, MEG [51], fMRI [52], fNIRS [53], etc. Overall, these techniques can enhance our understanding of the modular and specialized nature of the brain, causal order of temporal and spatial processing, and ubiquitous muscle functions, facilitating comprehensive analyses that advance neuroscience, psychology, medicine, and related fields. However, there is no known commensurate tool to measure similar features in deep networks (e.g. modularity) despite a few solid attempts that have been made in the information and statistics theory domain [54, 55]. The main objective of these mathematical tools is to primarily measure information modularity and evaluate robustness which results in leveraging algorithmic techniques to force models to demonstrate similar characteristics concerning processing specializations.
One of the most fascinating phases of human brain development is postnatal period, covering early few years of infancy where new brain connectivity suddenly transpires at insane rates in light of neural prolifiration and subsequent prunning processes [56]. This dynamic mission of architectural modifications can be shown to be beneficial even in normal brain development [57]. In the meantime, frontend optics as well as retinal anatomy also develop and start allowing more light to enter and higher number of cells to participate in retinal image processing, respectively, for more precise spiking, higher-level and accurate perception. It is shown in several past studies that poor visual acuity in human newborns offers several advantages, including faster and more accurate acquisition of visual object categories at multiple hierarchical levels [58]. These structural changes in every part of visual system such as chromatic and contrast sensitivities and its impact on the backend cortical substrate and functional connectivity can be used to design biologically-plausible (or so called biomimetic) training regimens to help train the machines in a similar way [58, 59, 60]. For instance, a training regimen that begins with high blur and gradually reduces its effect over time (allowing more high frequency to gradually enter), mimicking human development, is shown to enhance basic-level categorization performance compared to using non-blurred inputs throughout the training. These progressions may hold adaptive value by fostering the formation of receptive field structures that confer resilience to subsequent spatial or chromatic degradations. Fig. 4 demonstrates an example of biomimetic progression of acuity and color, applied to a toy image. As can be seen while the initial days of post natal period sees black and white, the maturation of cones leads to gradual processing of color later in development. However, in human development, not only are spatial frequencies and color subject to dynamic changes, but also the anatomical structure, generated internal noise over time, and maturation of recurrent pathways also contribute to this biomimetic progression. These factors collectively perfect the input feedback-consolidated architectural refinements over time, more so in the infancy [61, 62, 63].
One of the well-known barriers to good generalization is the overfitting phenomenon in machines. To practically combat against overfitting, regularization is proposed by injecting noise into the system to ensure bare minimum robustness [64, 65] or expedite the training time [66]. The addition of noise into training is quite similar in spirit to noisy optimizations within the context of simulated annealing [67, 68], which can help the model escape local minima by allowing it to explore a broader region in the loss landscape. By introducing noise, the optimization process can avoid getting trapped in small basins and suboptimal points, thereby improving the chances of finding a potential global minimum located inside a large basin. This mechanism is particularly useful in deep neural networks, where the loss landscape is highly non-convex with many local minima [69]. Understanding the learning mechanisms in humans, particularly the effects of attention and focus on the learning process, is crucial. These biological phenomena parallel computational techniques that mitigate erroneous learning and facilitate fact-checking, thereby ensuring the accuracy of information processed by the belief subsystem.
Whether developmental changes in neural noise adhere to a similar rationale researchers has found in acuity, contrast sensitivity and color changes is still open. An attempt has been made recently to explore the interplay of simultaneous progressions of acuity, color and noise [70]. These results suggested that it is not unreasonable to expect that the temporal progression in neural noise could contribute to improved categorization performance, particularly under challenging environmental conditions. The more intriguing question is the computational mechanisms that the brain puts in place such as generating the supply of noisy transmissions to realize this performance improvement. Unlike linear systems, which struggle to model complex behavioral dynamics and suffer from reduced information loss in the presence of noise, non-linear systems like the kind found in the brain can express multi-stability and explore various functional network configurations, provided sufficient internal noise is present [71]. Internal noise is an intrinsic part of neural dynamics, yet in many practical scenarios, it is also influenced by external stimulation and visual (perceptable) noise. This interaction is significant due to the emergent behaviors it produces, commonly referred to as stochastic resonance (SR) in biological systems. SR manifests when a nonlinear system’s signal-to-noise ratio is enhanced at moderate noise levels. At these noise intensities, the signal is able to reach the threshold without being overwhelmed by the noise, thereby optimizing signal transmission [72]. As a result, the interplay between non–linearity and internal noise becomes crucial for the spontaneous emergence of exploratory dynamics, even without external stimulation or noise. Observation of such biological phenomenon could be extremely useful and inspire us to design and train computational models that would have human–like behavioral dynamics in resolving complex tasks.
Yet another source of inspiration could relate to the specialized brain pathways to process distinct spatial frequencies for enhanced task completions under severe environmental conditions. Early studies on spatial frequency disentanglement in the brain suggest that low and high frequencies are processed distinctly, involving separate visual pathways that develop later in life to enhance human robustness in identification and categorization tasks [73]. The most straightforward translation of human studies to machine model development or learning involved altering training regimens to include statistically relevant data or introducing various degradations in the image space to align the underlying distributions driving training and testing conditions. However, biological processes follow a path coupled with the environment where external stimuli directly affect the formation of new neuronal pathways, resulting in the proliferation or pruning of neural connections and the formation of specialized neural populations that fire together. Simultaneous firing eventually forms strongly connected networks (wiring in Hebbian sense) in such specialized brain regions. Having observed such biological mechanisms naturally encourages researchers to investigate the necessary architectural changes required for machines to acquire human-like robustness, which can be provided by, for instance, the modularity of brain functionality.
3.2.2 Human–assisted AI: Hybrid Systems
In Human-assisted AI, models are developed with the primary goal of augmenting or enhancing human/machine capabilities rather than replicating machine or human intelligence. Human-assisted AI leverages AI technologies to assist machines or humans in performing tasks more efficiently, accurately, or autonomously. Human–assisted AI, as a bidirectional enterprise, can be characterized as the continuous interaction between human-helping-machines and machines-helping-humans. However, fully realizing the potential of human-AI collaboration requires to address important challenges. First, the necessary conditions that support complementarity – where human-assisted AI performance is expected to surpass that of AI alone for a given task, requiring humans to recognize when to augment AI that complement critical decision-making processes. Second, accurately assessing human mental models as well as the extent to which humans are prone to failures are crucial, which also brings out the issue of understanding the effects of design choices in human-AI interaction, such as the timing and amount of Human assistance to avoid bias.
Primary example of this trend is the foundation models that can interact with humans, such as most recent advanced LLMs [74] (GPT4 [6], Llama [75], etc.) that begin transforming the software industry by providing assistance writing the bulk of the required code as a downstream task, and still subject to thorough validation [76, 77]. This approach (especially the former reference [76]) emphasizes the significance of collaboration between humans and machines, with AI systems providing support, guidance, or automation to complement human skills (critical thinking, reasoning, navigation of complex decision processes) and expertise. Other examples include AI-augmented tools for object recognition, memory enhancement [78], data analysis, medical diagnosis [79], language translation, and robotic assistance in manufacturing or logistics [80]. Yet another hot topic could be mission critical applications where humans and machines could perform tasks using different strategies and hence end up having distinct failure patterns. This complementarity can be utilized as a fusion to bolster overall accuracy performance [81].
Human mind organization and sensorimotor systems could be useful for developing computational systems in various ways. One remarkable example from vision is the leverage of eye tracking and analysis of gaze direction, which has the potential to uncover a wealth of information about internal cognitive states. For example, it is argued that gaze can reveal the underlying attentional patterns [82]. More broadly, gaze information could be useful from two standpoints. It can help AI develop an attention mechanism to effectively cope with the information-rich world and determine significant cognitive features [83]. In fact, given the amount of training data required to learn such an adaptive attention process, it may be easier for AI agents to learn attention directly from human gaze data. Secondly, it can help AI interfaces to perceive and understand gaze movements to create naturalistic interactions with humans [84, 85]. For instance, gaze-based control mechanisms could allow users to interact with AI systems without physical input devices (keyboards, mouse, etc.), fostering a more immersive and responsive interaction.
In summary, while human-inspired AI seeks to replicate human intelligence, human-assisted AI focuses on enhancing human capabilities through collaboration with AI technologies. As collaboration and interaction between these two forms of intelligence occur, both will inevitably influence each other, leading to potential evolution and changes in their characteristics over time. However, both approaches originally have distinct goals and applications, and the choice between them would depend on the specific needs, objectives, and ethical considerations of AI development and deployment.
3.2.3 Human–independent AI: The rise of Open Intelligence
In our taxonomy, this branch is still considered as human-centered AI because the studies in this category are still framed based on human intelligence as a reference model. For instance, although AlexNet [34] is inspired by the visual cortex and its layered functional structure, designed to mimic human visual perception and processing, the proposed optimization algorithm backpropagation (also referred as back-prop) is not necessarily human-inspired; it is more inspired by how computational constraints shape and efficiently utilize resources, along with mathematical convenience such as sticking to the same set of weights/coefficients in both forward and backward passes [86]. In fact, back-prop is considered to be an optimization method which spreads information across all parameters of the neural network making the network less modular [55] (also see Section 4.3). Hence one can argue that the development of AlexNet is not necessarily categorized as human-inspired but rather neuroscience-inspired at the very high-levels of abstraction subject to SoTA silicon hardware constraints and explainability.
The main motivation behind Human-independent AI is twofold: (1) In pursuit of an open intelligence that is independent of human intelligence could potentially be a better model for the natural world we live in as human intelligence seems to have missing elements to tackle important issues of the world such as less advanced predictive capabilities for extended time and range or the inherent limitations of convolution/pooling operations executed by the visual system. This is corroborated by the most recent transformer models which lifts off some of the important inductive biases that previous generation CNNs adopted. Second observation to this idea is the ever growing data scales as machines having less inductive bias would require to build such biases based on the voluminous data. (2) In any reliable/efficient AI system development, we would expect a joint symbiotic life between the intelligence and the hardware that runs it. The current silicon hardware and Von-Neuman architecture is not necessarily the right home for the optimal implementation of human intelligence. This might guide us to think of other forms of intelligences that could optimally run on the available hardware architectures, building on the few-decade experience of distributed microprocessor history of human construct. For more elaborate discussion on the effect of hardware on the implemented intelligence, please see 5.3.1.
Perhaps it might be more accurate to frame human-independent AI as an effort which characterizes humans as irrelevant to the next generation large-scale AI development. There is a growing inclination in the community believing that the independent design will benefit from some of the findings that come out through interactions with the outside world (experimentations etc.) which might well be counter-intuitive and cannot be adequately represented by the data intrinsics. However, these designs are still bound to subjective evaluations by human-credited metrics and might be compared to human performance to improve their explainability. After all, explainability is a human-construct and only meaningful if it helps with understanding the true nature of phenomenon.
Figure 5: Elements of Brain-inspired information processing.
4 Brain-inspired Information processing
Brain-Inspired Information Processing refers to building computational systems and designing architectures that emulate the brain’s structure and functions to achieve advanced cognitive and data processing abilities. This interdisciplinary field draws on neuroscience, cognitive science, and artificial intelligence to develop models and systems that implement brain-like processes for tasks like multi-modal perception, learning in development, and critical decision-making/building commonsense. The overall picture is illustrated in Fig. 5 where the elements of brain-inspired information processing as well as the context is introduced in which this paradigm is utilized [87]. To get a grasp on how the brain operates, various neuroimaging techniques are utilized, including but not limited to Electroencephalography (EEG) and Magnetoencephalography (MEG) which measure the brain’s electrical and magnetic activity, offering high temporal resolution for tracking dynamic neural activity [88]. Functional Magnetic Resonance Imaging (fMRI), on the other hand, provides detailed spatial resolution by monitoring blood flow changes, helping to map which brain regions are involved in particular cognitive functions [89]. These tools are crucial for revealing the neural dynamics and connections that underpin complex cognitive processes. To identify neural signatures for different brain functions, machine learning algorithms are trained on brain signal patterns associated with specific tasks or mental states. By analyzing the relationships between brain regions, statistical methods (such as random graph theory) help model brain networks and pinpoint crucial regions/nodes involved in processing information [90].
The informative features extracted from the biomedical data are typically used to build computational models that mimic the way our brains process information. Translating these measurements into computational frameworks that are truly brain-inspired could be quite challenging though. For example, neural networks can be structured to reflect the hierarchical nature of the brain, with different layers corresponding to various levels of processing found in, say, the visual cortex [91, 92, 89]. Finally, these computational frameworks are refined by comparing them against real brain data, perhaps at higher abstraction levels, ensuring they not only mimic brain activity but also can generalize to new, unseen tasks. However, the resemblance is always subject to questions as it is hard to completely translate brain operations to algorithms and implement them on silicon. By blending neuroimaging with AI and statistical analysis, researchers mainly aim to create brain-inspired models that capture, at least to some extent, the core manifestation of human cognition. These models can then be applied to fields ranging from the most popular foundation models such as LLMs to autonomous systems. In fact, brain-inspired data processing could guide foundation models to embrace emergent behaviours more rapidly as architectural similarity is more likely to demonstrate functional similarities [93] (See also Section 7 for further discussions). This approach seeks to bridge the gap between biological intelligence and machine intelligence, pushing toward more flexible and robust AI systems.
To fully adapt brain-inspired processing, it is crucial to tackle several fundamental questions, which can be grouped, to our count, into four key areas: (1) the possibility of reverse engineering of human cognitive abilities, (2) the distinctive features of the human brain, (3) the mechanisms through which our brains achieve modularity and robustness, and (4) exploration of gaps in current state-of-the-art approaches that prevent us from reaching human-level AI capabilities.
4.1 Reverse engineering a human skill
One of the main pillars of brain-inspired research is to understand how hard it is to fully grasp the various emergent behaviors that leads to a form of intelligence in the brain. This question dates back to 1980s and the difficulty of adapting it depends on the developmental trajectory of that skill. The premise of the past work was that we typically expect the difficulty of reverse-engineering any human skill to be roughly proportional to the amount of time that skill has been evolving in living beings including animals [94]. For example one of the oldest human skills is perception which is largely unconscious and so appear to us to be effortless. Examples are abundant: Anything to do with attention, visualization, motor skills, tactile system, social skills and so on. The hypothesis is that we should expect skills that appear effortless to be difficult to reverse-engineer. The dual to this hypothesis is skills that require effort may not necessarily be difficult to engineer at all.
Emerging skills developed later in the evolutionary trajectory such as mathematics, engineering, gaming, logic, and scientific reasoning pose challenges as they diverge from our evolutionary predispositions. These abilities have developed relatively recently in historical terms, undergoing refinement mainly through cultural evolution over the past few thousand years. Initially, many prominent researchers assumed that once they had grappled with the ”hard” problems, the seemingly ”easy” challenges of vision and commonsense reasoning would fall in place. However, the last decade in AI has proven it wrong. This misconception stemmed from the realization that these supposedly straightforward problems are, in fact, incredibly difficult, leading to the machines’ proficiency in solving problems like logic and algebra was irrelevant, as these tasks are exceptionally easy for them to handle [95, 16, 96].
4.2 What makes humans special and/or not so special?
Humans exhibit unique strengths and limitations in tasks like recognition, perception, detection, and search, which are shaped by both our biological makeup and cognitive abilities. First and foremost, humans are not adept at “detection” tasks like blur identification or identifying misaligned lines. Especially without focused attention, their performance significantly degrades, making it difficult to excel in completing such tasks successfully. Additionally, humans struggle with ”hidden correlation” tasks, as their ability to estimate the joint probability distribution of a large set of random events is very limited [97]. This limitation arises from the cognitive load and complexity involved in processing and integrating multi-modal sources of imprecise information simultaneously.
However, human cognitive system is extremely good at trying/searching (visual search) [98] and finding out things [^3] We will define “things” later in the context, predominantly referring to objects, entities, etc. (in other words, quite good at ontologies – a set of concepts and categories in a subject area or domain that shows their properties and the relations between them) leading to good performance in languages. Humans also have a sparse but adequate understanding of physical reality and causality which leads to data processing efficiencies in their generalization capacities. Due to this valuable feature, their learned world-models usually turn out to be reasonable and in congruent with the required tasks. They are also good at grounding the language with mental simulation. It is widely recognized that the way humans perceive the world is organized around objects - mostly up to 3rd dimension and some lausy understanding of the 4th dimension. These objects act as fundamental building blocks for a variety of higher-level cognitive processes, including language, planning, and reasoning. The notion of the world being comprised of separate parts that can be processed autonomously and combined in countless ways enables humans to make generalizations that go well beyond their direct experiences.
In the context of mental simulation, the ladder of causation is pertinent to our discussion [99, 100]. We should consider our question with respect to the three rungs of the ladder: (1) The initial stage involves the extraction of meaningful correlations from real-world representations through a process of observation. This endeavor aims to estimate the conditional expectations of events, thereby facilitating the formation of crucial associations. (2) The subsequent stage entails engaging with the environment using an accurate model of the world which might be formed partially in the previous rung. Interactions in this phase result in alterations to the characteristics of objects that are inherently linked or correlated, either directly or through other dependent or confounding variables. Given that such interventions occur over time, they engender a sequence of actions and reactions conducive to causal reasoning. (3) The final stage of the ladder, termed counterfactual thinking, necessitates envisioning alternative scenarios, evaluating them, and adjusting accordingly. This aspect is distinguished by its requirement for imagining changes and their potential repercussions, without necessarily engaging in active intervention. Achieving counterfactuality underscores one of humanity’s most demanding cognitive abilities: predictive reasoning.
4.3 A case for Modularity and Robustness
The brain network modularity can be defined as the degree to which functional brain networks (connectivity) are divided into special subnetworks. Modularity is assumed to be beneficial to develop a robust brain against internal as well as external perturbations. Biological individual differences seem to suggest the mandatory modular organizations for the robustness of persistent activity to various perturbations [101]. Modularity does not typically form in computational models with standard training regimens and architectures. However, modularity can be enforced through various approaches such as changing the training or the optimization process. Biomimetic training for instance is shown to demonstrate more robust behaviour under out-of-distribution scenarios. In addition, more biologically plausible optimizations provide robustness and seem to allow the emergence of invariant representations more quickly [102].
One way to achieve robustness in recognition tasks is to retain information in redundant and distributed manner. A widely accepted assumption is that distributed information leads to robustness to internal and external perturbations, whereby lost or corrupt information in one brain region can be compensated for by redundant information in other regions [103]. Redundant modular representations might naturally emerge in neural network models that could learn robust dynamics [101] as long as such representations can be retained in later stages of training. Indeed, developing modularity might allow learned features to be frozen and the rest of the other hardware could be used to learn new tasks allowing learning rates to change over time for different parameters of the neural system [104, 105]. As can be seen, modularity has the potential to lead a system to be more robust to external perturbations. However, not every robust system necessarily manifests any sort of modularity in their operational characteristics.
4.4 Missing elements towards achieving Human-level AI
Here is a corollary that could well easily be established through a rigorous search on the AI research in the last decade or so. It summarizes the main elements missing in any design towards achiving human-level capabilities [95, 106, 27].
Without higher level abstractions, counterfactuality, predictive reasoning and explicit compositionality, it is almost impossible to achieve or get to human-level AI in the near future.
Tools for few-shot generalization (without exponentially increasing the data size for training) [107], must be totally different than what AI community is allocating their attention on/and may require the next breakthrough in science and engineering. One approach to achieve such immediate generalization could be “biological plausibility” [108, 109] which is often exercised by the research community in diverse ways including the modifications of neuronal structure-archıtecture [109], inductive bias, training regimen [110], activation patterns etc.
Biological plausibility in neural network design and training refers to how closely artificial neural networks (ANNs) mimic the structure and function of biological neural systems, like spiking neurons in the human brain [111]. This involves replicating neural architecture, using local learning rules akin to Hebbian learning [112] or even the advanced BCM theory [113], incorporating modification thresholds, learnable activation functions and dynamics, similar to biological neurons, achieving energy efficiency and provide plausible explanation for synaptic scaling [114]. However, biological plausibility must be understood in the context of the scale at which learning is considered. For instance, if a deep network is viewed as an unrolled recurrence in the brain, an argument can be made in favor of its biological plausibility [15]. Conversely, when examining fundamental operations such as convolutions and back-propagation for parameter optimization, an argument can easily be made against its biological plausibility [115, 86].
Networks that are biologically plausible can improve interpretability, robustness, adaptability, and energy efficiency, reflecting some of the advantageous properties of biological systems. However, biological plausibility does not necessarily ensure behavioral similarity, which is the resemblance in outputs and responses to those of biological systems. Achieving behavioral similarity requires addressing additional factors such as the complexity of tasks, generalization capabilities, and learning efficiency. While biologically plausible networks may show some behavioral similarities due to their design, achieving full behavioral mimicry involves broader considerations of intelligence, learning, and adaptability beyond just structural and functional resemblances.
Another missing item could be the compositional understanding of symbol-like entities such as objects. This encompasses the ability to recognize and represent objects not merely as isolated entities but as components of larger, more complex structures. In biological systems, this compositional understanding allows for the hierarchical processing of sensory information [116], enabling the perception of scenes, the understanding of contextual information, and the recognition of relationships between objects and their affordances. Incorporating this level of compositionality into ANNs could significantly enhance their ability to generalize from limited data, perform complex reasoning tasks, and interact more effectively with dynamic and multifaceted environments. This involves developing models that can dynamically form and utilize higher-order representations, mimicking the way the human brain integrates individual symbols into coherent and meaningful wholes. One particular computational manifestation of increased compositional understanding could be attributed to effective receptive fields of the stacked layers in CNNs [117], which could potentially be achieved either through using larger kernel sizes in convolutional layers [118] or deeper architectures with smaller kernel sizes (such as ResNet etc.).
It is also subject to question to reevaluate the biological plausibility of basic operations of deep networks such as convolutions, activation functions or optimizations (learning rules) etc. One of the most critical aspect of convolutions used in computer vision is the weight sharing. State-of-the-art convolutional networks utilize shared weights, a capability that biological networks, which require local weight updates, do not possess [119]. Several studies have attempted to relax weight sharing in convolutional networks by introducing locally connected networks [120, 121]. However, such networks perform worse for large scale architectures and hard tasks e.g. large-class categorization.
Another most significant barrier to achieving human-level AI might be the way the learning is interpreted and implemented in neural network models especially in the most recent foundation models [122]. The principles underlying backpropagation can inspire the development of biologically plausible learning algorithms. Mechanisms such as feedback alignment, burst propagation, and predictive coding offer promising alternative pathways for approximating the learning capabilities observed in ANNs. The work by Lillicrap et al. [123] highlights the importance of these alternative mechanisms, bridging the gap between computational models and biological reality. While backpropagation itself may not be biologically plausible, its core concepts can be used to guide the learning at an higher level and continue to influence the understanding and modeling of learning in the brain.
5 Challenges and Perspectives in Human-Level AI Development
5.1 Challenges in the development of human-level AI
There are many challenging and often obscure obstacles in the development of human-level AI. Addressing these issues requires interdisciplinary collaboration across diverse fields such as computer science, neuroscience, psychology, philosophy, and ethics. This multifaceted endeavor comes with perhaps the most intriguing unsolved problems of science and engineering.
One of the most fundamental issues in the deveopment of human-level AI is concerned with the identification of mechanisms that lead to commonsense reasoning in complex systems such as humans. For one thing, AI systems that possess a reasonable understanding of commonsense knowledge are expected to make sense of everyday situations and make intelligent decisions based on implicit knowledge that humans take for granted. As indicated in [27], common sense in AI are expected to build representations of core physical constructs such as time, space, causality, and establish interactions between physical objects and humans. Through the leverage of abstraction and compositionality, they should support reasoning techniques capable of handling complex and incomplete knowledge. More fundamentally, they also need to possess a human-inspired learning system (also see section 4.4) that continually incorporates new knowledge from various interactions and sources, enabling deep understanding and general intelligence. While carefully putting together bits and pieces, structural resemblences to human anatomy should also bring together explainability and interpretability components, which should naturally fall in place due to the joint consideration of learning and architecture. Creating AI that can explain its decisions and actions in a human-understandable manner is very important for building trust between machines and humans. The mutual trust is especially important in mission-critical applications such as healthcare and autonomous vehicles, where understanding the rationale behind AI-driven decisions can be a matter of life and death.
AI systems must hold up ethical principles and societal values as part of their design and ensure that algorithms running on them adopt fair, transparent, and unbiased behavioral characteristics. This undertaking involves navigating issues such as privacy, accountability, and the long-term impact of AI on biological intelligence as well as employment and other dynamics of society at large. It is not a surprise that the next generation AI and its implementation will generate significant societal impacts, necessitating moderation and legislative regulations.
Training foundation AI models and enhancing their ability to generalize knowledge across different domains and tasks through specialized fine tunings, also known as transfer learning, is another major challenge. This adaptation to novel circumstances involves enabling foundational AI systems to learn novel concepts from only a few samples of data (a.k.a. few-shot learning [124, 107]), leveraging previously acquired knowledge to solve new problems more efficiently. However, as is the case with modeling many cognitive functions of the brain with regards to real-world problem solving, there is no widely accepted mathematical model of transfer learning that connects out-of-distribution generalization to the minimum number of model parameters required for updates. Despite easy to articulate, accomplishing such a goal involves many genuine techniques that would make machines look like humans. Of such mechanisms, meta-learning [125] and self-supervised learning [126] are the two other critical areas of human-inspired research. Developing AI systems that can learn to learn, adapt, and improve over time through meta-learning techniques are desirable properties of the future generations of this technology. The overall goal of learning involves designing algorithms that can automatically discover effective strategies and apply them to new tasks, allowing AI to become more autonomous, capable and self-directed over time.
Integrating AI systems with physical bodies (embodiment) and sensory capabilities to enable interaction with the physical world is another significant challenge towards human-level AI. This concept, known as embodied intelligence [127], involves understanding how intelligent behavior emerges from the complex interplay between an agent’s hardware components, implementation of the software logic and its dynamic environment. Achieving this integration is crucial for performing complex tasks requiring learning based on sensorimotor skills [128].
As the AI systems begin to replace many critical positions of human occupancy in making mission-critical and life-changing decisions, ensuring that they are safe, secure, and robust against adversarial attacks, system failures, and unintended intrusions gradually become a major focus of today’s research. This includes developing techniques for verifying and validating AI behavior and mitigating potential risks associated with their deployment. Maybe the missing elements could be as far-stretched as developing self-awareness and characterization of self-assessments. The conceptualization of consciousness and self-awareness in AI systems remains one of the most widely studied and elusive goals [129, 130]. Investigating the nature of consciousness and exploring whether it is possible to create artificial entities that possess subjective experiences such as feeling secure or safe, similar to humans is still largely unknown. Although there have been attempts at mathematical modeling [131, 132], these have not yet sufficiently explained the complete conscious experience that humans present.
5.2 Transferability of human-specific attributes to machines
In order to transfer a human specific attribute (Ethics, Aesthetics, Coinciousness, Morality, Humor, Causality or Trust), it is a priority to define these attributes using a precise language. Without a clear systematic description, it would be hard to comprehend and even have a common consensus on their meanings. However, creating a clear understanding entails improving the precision of their dictionary definitions [133], which might not be an easy task to perform in a mathematical sense. In fact, having precise definition for biological and metabolic events could sometimes be pretty hard to frame, which are analyzed historically using a branch of mathematics known as information theory [134]. In fact, imprecision in the development of human attributes could be beneficial in accurate and robust decision making processes.
5.2.1 The role of imprecise nature of perception in learning
Human beings have evolved to rely on approximate reasoning as a fundamental strategy for navigating the inherent uncertainties of the world such as found in volatile environments [135]. This capability allows them to process and respond to imprecise, incomplete, or ambiguous information with remarkable flexibility and adaptability. For instance, when interpreting social cues, making ethical judgments, or appreciating humor, humans often operate within a realm of subjective interpretation rather than absolute precision. These traits—ethics, aesthetics, humor, and beyond—are not merely ancillary to human cognition but are deeply intertwined with the very nature of how humans perceive and interact with the world [135, 136].
Given this, there is a compelling rationale for incorporating the modeling and processing of imprecise information into learning and expert systems. By doing so, we can better emulate the complex, layered, and often ambiguous ways in which human reasoning operates. For example, in the development of AI systems designed to interact with humans, such as virtual assistants or social robots, embedding a capacity for handling imprecise information could enhance the system’s ability to understand and respond to human emotions, make ethical decisions, or even generate humor in ways that feel more natural and relatable.
Moreover, embracing the nuances of imprecision in AI models aligns with a broader movement in cognitive science that recognizes the value of “fuzzy” logic and probabilistic reasoning as closer approximations to human thought processes [137, 138]. This approach not only makes AI systems more robust in dealing with real-world complexities but also opens new avenues for creating machines that can truly resonate with the multifaceted nature of human experience such as multi-modality model development and multitasking. In summary, by modeling imprecision, we do not just aim to replicate human reasoning but to enrich AI’s capacity to engage with the subtle and sophisticated dimensions of 3D world model and human experience roaming within.
5.2.2 The effect of neuroscience studies for real-world stimuli
In traditional neuroscience research, most experiments are conducted within the confines of highly controlled laboratory environments. While these controlled settings are invaluable for isolating specific variables, formulating hypotheses and understanding fundamental neural mechanisms, they often fall short of capturing the complexity and variability of real-world experiences. To bridge this gap, it is increasingly recognized that incorporating real-world stimuli into experimental designs—integrating multisensory signals and introducing various confounds—can significantly enhance the ecological validity of neuroscience studies [139, 140].
By designing experiments that more closely mirror the diverse and dynamic conditions encountered in everyday life, researchers can gain a deeper understanding of how the brain processes the rich tapestry of sensory information received on a daily basis. For example, rather than presenting isolated visual or auditory cues, skilfully integrated into experiments could involve scenarios where participants navigate through virtual or augmented reality environments, responding to a combination of sights, sounds, and even tactile feedback in different settings such as brain-computer interfaces [141, 142]. This approach not only provides insights into how different sensory modalities interact in the brain but also reveals how contextual factors, such as distractions or emotional states, influence perception and decision-making.
Furthermore, studying the brain’s response to real-world stimuli can uncover how neural processes adapt to complex and unpredictable environments. This line of research is particularly relevant for understanding higher-order cognitive functions such as attention, memory, and problem-solving, which are constantly challenged by the multifaceted nature of real-world tasks. By moving beyond the artificial simplicity of lab-based experiments, neuroscience research can begin to unravel the brain’s remarkable ability to filter, integrate, and make sense of the overwhelming amount of information, encountered daily in natural settings.
Ultimately, this natural shift towards more realistic experimental paradigms holds the potential to transform our basic understanding of brain function, offering new insights into how neural circuits support adaptive behavior in the real world where the brain is constantly exposed to multi-modal sensory input. Such advances could pave the way for more effective interventions and technologies designed to support cognitive function in everyday life, from improving educational tools to developing brain-machine interfaces that operate seamlessly within the complexities of the real world [143, 144, 141].
5.3 Alignment of SoTA approaches with human-inspired/human-assisted AI
There is no doubt that today’s AI models are useful in closed and regulated application domains addressing successfully the specific task requirements. There are various angles of criticism directed at these systems, including algorithmic bias, ethical concerns, safety issues, overreliance on their use, ethics and their impact on creativity development, among others [145]. Rather than diving into the details of each critical point, which is what past research have heavily done so far, I will rather explore the alignment of current models in the context of hardware, architecture built on this hardware, and the learning schema that generates the software in the following subsections. We shall start considering the effect of the today’s hardware architectures (Von Neumann or parallel architectures of GPUs etc.) on the trajectory of AI system development.
5.3.1 The influence of the architecture of today’s computational models
Let us begin addressing this question in the context of Marr’s three levels which states that any complex information processing machine must be interpreted through the following three levels of organization [110].
  • Computational theory: Goal and appropriateness of the computational logic or strategy.
  • Representation and algorithm: Representation for the input/output and how the algorithm transform one to the other.
  • Hardware and implementation: The physical implementation of the algorithm.
Marr’s approach argues that complex systems have multiple levels of organization, forming a hierarchy where each level consists of composites of the entities at the lower level. These systems can be interpreted at various levels, such as biochemical, cellular, and psychological levels in organisms or logical gates, machine language, and high-level tasks in computers. Higher-level explanations allow for generalizations across different underlying physical structures. This means we can predict behaviors without detailed knowledge of the underlying components, like predicting genetic traits without knowing DNA specifics or predicting a computer’s output without understanding physical properties of semiconductors. The power of higher-level explanations lies in their ability to offer reasonable predictions with less detailed information about the system behavioral details. They achieve generality by focusing on the input/output functions of subsystems rather than the functional significance of their inner workings.
The success of today’s AI heavily dwells on the data representations and algorithm developments just like we see it in the development of neural networks and the algorithms that enabled such computational models become realistic such as Bayes updates, gradient descent, dynamic and probabilistic methods and theories. The main medium used to implement AI is silicon-based transistor logic employing Von Neumann architectures, characterized by large, inactive memory and minimal state within the central processing unit. Due to the significant bottleneck posed by single processors and low-latency local memory in meeting the demands for real-life workloads and pervasive bandwidth demands, parallel architectures have garnered importance, enabling independent segments of programs to execute concurrently. Consequently, numerous AI model development endeavors appear to be gravitating towards maximizing the efficiency of parallel architectures for training. A notable instance of this trend is the recent advancement in transformer models, which exhibit a high degree of parallelizability compared to convolutional neural networks [146, 147].Of course such a bias leads to a broad range of changes in the way the algorithms are designed, ultimately changing the implementation details of intelligence on the silicon.
Decoupling the implementation from the set of required computations for, say, visual perception should enable the development of human-like AI on silicon substrates without any fundamental limitations. However, this assumption hinges on a complete understanding of the computations underlying visual perception. In other words, once the functionally realizable computational steps governing perception are fully elucidated, we should be able to implement these steps with sufficient instructions to achieve high-level human-like interpretations. Nevertheless, the biological computation framework remains unresolved, and we frequently rely on computational models to understand complex biological systems. Hence, our interpretations of biological hardware and how learning transpires become unequivocally dependent on the way hardware is manufactured, including the architecture, power use and data handling.
One key distinction between biological and digital intelligence is the strength of the coupling between the software implementing the aspects of intelligence and the available hardware this software is designed and built on. Abstraction between the layers in any complex system development allows modularity between system components, making design manageable but complicating debugging and failure tracing. Additionally, as the data volume increases and the models get chunkier, the implementation of digital intelligence seems to require significant resources, exponentially scaling the required power to support their learning. However, digital intelligence, with its parameter values, network links, and overall model descriptions, is highly efficient for knowledge storage and transfer across different digital hardware, reflecting its decoupled design nature. In otherwords, the learning process, despite being heavy and overwhelming, could be executed only once for all and then transferred/distributed for further learning through specializations, which also highlights the significance of foundation models. In contrast, human intelligence is tightly coupled with biological hardware, offering physiological advantages besides its efficient analog knowledge representation and processing. This tight coupling makes the overall learning process quite challenging, as each hardware needs to be trained from scratch using innate knowledge integrated into the architecture.
5.3.2 Model Architecture and Learning
Model Architecture
In the evolving landscape of modern machine learning models in computer vision, Convolutional Neural Networks (CNNs) and the Vision Transformer (ViT) stand out as two of the most impactful approaches, each offering distinct architectural and representational features [148]. While CNNs have long dominated tasks like object detection and image classification, recent advancements in (multi-head) attention-based Transformer architectures, particularly ViT [149], challenge the conventional reliance on convolutions in computer vision. Despite the recent work [150] showing the expressive equivalence of naive convolutions and the implementation of attention mechanisms, the advent of transformer’s internal components points at alternative feature extraction techniques which may or may not align with brain’s intricate processing of objects (which can be argued to be akin to mechanistics of convolution operation).
Convolutional Neural Networks (CNNs) have been the de facto standard for a range of computer vision tasks, and significant efforts have gone into understanding their learned representations on popular datasets like ImageNet [9]. The emergence of diagnostic datasets, such as Stylized ImageNet [151] or CIFAR10-c and CIFAR100-c [152], has deepened our comprehension of model biases. For example it is shown that most of the pre-trained large scale CNN models are biased towards texture rather than shape unlike human categorization processes [151, 153]. As is the case with the human vision, CNN’s convolutions are adept at capturing local patterns in images and exhibit translation invariance, making them robust to variations in position and orientation of objects within the image. In addition, CNN’s layered architecture can be considered more interpretable with respect to functional organization of human visual system, as features learned in earlier layers often correspond to simple patterns like edges and textures, while higher layers capture more complex features [154]. As a computational feature, parameter (weight) sharing in convolutional layers help them be more memory efficient. However optimizing parameters for the entire image space can be seen as quite restrictive not to mention the fixed input size. The static nature of the layered structure as well as the optimization (learning) are nowhere near how the brain operate, and are structured and trained.
However, the landscape has evolved with the introduction of various tweaks for Transformer architectures like Vision Transformer (ViT) [149, 155], known for their flexibility and success in time-series data analysis and natural language processing tasks. They now form the fundamentals of foundation models [122] with large scale distributed hardware support. ViTs capture global dependencies in images by treating them as sequences of patches, allowing them to potentially model long-range interactions more effectively. Such patch-by-patch processing could be quite scalable as high-resolution images can easily be learned without multiple runs of tedious processing and plus workload independence lends this processing to be easily executed on parallel architectures.
Recent models have key components that make them unique in the learning trajectory such as Convolution vs. Attention: While CNNs demonstrate power and scalability, their reliance on local connectivity raises concerns about global context loss. Transformer models, rooted in attention mechanisms, provide an alternative that challenges this paradigm. The introduction of Transformer architectures to computer vision adds a layer of complexity and adaptability, assisting vision models to infer long range dependencies more effectively. Superiority of Transformers is largely credited to their self-attention-like architecture which might not necessarily be biologically plausible. There have been many attempts to improve the effective receptive field size of these networks to compensate for the long range dependencies [118].In fact, in [156], authors highlights three highly effective architecture designs for boosting robustness: (1) patchifying input images, (2) leverage of larger kernels and (3) reducing activation layers and normalization layers. Although attention-based activations can be shown to be more generalized operations and can implement convolution [157], their presence seems to be not mandatory to ensure robustness in various tasks of categorization and recognition. These findings indicate that there are likely several architectural options available to implement specific biologically plausible functionalities, thereby ensuring robustness. Indeed, there are already few published works arguing about biological plausibility i.e., neurons can perform attention-like computations through short-term, Hebbian synaptic potentiation [158]. When an axon’s activity matches a neuron’s activity, the synapse is briefly potentiated, enabling the axon to control the downstream neuron.
Model Learning
Biological and computational learning are mechanistically different. While the biological system strengthen or weaken neural connections based on experience, primarily through mechanisms like synaptic plasticity, computational systems simulate learning through scalar/vector weight values optimized by data, optimization and appropriate use of cost functions. Neither the data nor the optimization are necessarily bio-inspired. Among many other distinctions between biological and computational systems, optimization and locality of learning rates are two distinctive features of mechanisms employed in the learning process.
Evolution as a computational approach to biological optimization refers to the application of evolutionary algorithms to solve complex and often constrained optimization problems. These algorithms mimic the mechanisms of mutation, crossover, and selection to iteratively improve solutions, leveraging biological principles to find optimal or near-optimal configurations in various domains, including genetics, neural networks, and ecological modeling. This approach capitalizes on the inherent efficiency and adaptability of evolutionary processes to address intricate challenges in biochemical mechanisms and computational biology. Although backpropagation is not necessarily required for networks to attain biologically interesting properties [159], it might be a good idea to leverage the main properties of ecosystems in building computational tools to solve complex problems [160, 159]. The possibility of inserting ecological concepts in the optimisation process is key in the development of novel bio-plausible hybrid systems.
One of the fundamental issues of computational intelligence is the non-local learning rules, which poses a challenge because it remains unclear how individual neurons and synapses in biological circuits access information about the global state of a circuit or a specific optimization target. The sources highlight the ongoing research focused on translating these global rules into local synaptic learning rules that rely solely on information locally available to neurons and synapses, such as pre- and post-synaptic spike timing [161]. Related to this regional learning, one of the stregths of biological networks is that they typically balance learning among multiple competing objectives, whereas computational schemes focus on optimizing for a single task. Future research could explore incorporating factors like energy consumption, wiring length, processing speed, learning efficiency, and robustness into ANN optimization to gain a more comprehensive understanding of biological circuit design.
5.3.3 Comments on building blocks: MLP, KAN and beyond
The Multi-Layer Perceptron (MLP) [162] is a widely adopted model to simulate biological information processing through continuous approximation mechanisms. In other words, it is a crude representation of high-level functionalities of synapses and axons/dendrites in the human brain, making it easier to apply calculus rules and to implement learning on digital hardware with the available instruction sets. The generalized use cases of MLP consists of multiple layers of interconnected nodes, or so called neurons, where each neuron performs a weighted linear sum of its inputs followed by a non-linear activation function. Due to Universal Approximation Theorem (UAT) [163], this intuitive structure enables MLPs to approximate any arbitrary non-linear continuous function by learning and modeling complex patterns and relationships within data.
Arnold-Kolmogorov (AK) theorem suggests that any multivariate function can be decomposed into a sum of compositions of univariate functions, offering a theoretical basis for designing networks known as Kolmogorov-Arnold Networks (KANs) [164]. These networks leverage the theorem to simplify the construction of neural networks capable of approximating complex functions with fewer parameters and potentially more efficient architectures in which weightening is replaced with individual spline-based activation functions.
However, we need to realize that the UAT and the AK theorems are two alternatives for multivariate complex function approximations and have found distinct implementation styles, inspiring computational hardware designs that might emulate biological information processing better. Concordantly, MLPs and KANs represent two different but similar approaches to neural network architecture design as well. While MLPs focus on signal linearity, flexibility, and generalization capabilities with voluminous data, Kolmogorov-Arnold Networks aim at non-linearity at the edges, optimizing the representation and computational efficiency. Despite the potential of reducing the number of parameters, theoretically, we can make a few observations for these networks: (1) MLPs with learnable activation functions that could be formulated in terms of the sum of splines have the potential to approximate activations of KANs. (2) This theoretical equivalence leads to the conclusion that the main issues with MLP-based network design such as the curse of dimensionality are still applicable to the fundamentals of KANs.
On the other hand, theoretical equivalence does not necessarily mean the practical implementation of these networks is the same in terms of power usage, parameter optimizations, training time, hardware amenability, complexity of implementation, etc. For instance, let us think about hardware implementations optimized to perform spline activations or spline parameter optimizations, and hence while the hardware deals with the non-linearity, the software can run higher-level simpler operations such as linear sums, thus enabling complex operations to run closer to digital hardware through creating close couplings. These instruction-level couplings could be a source of further optimizations in terms of power consumption, transistor counts per processed bits or processing latency etc.
5.3.4 Texture/Shape Bias and Fine-Tuning in vision models:
In a relentless pursuit of building models that better align with human perception, researchers have been focusing on two key features: The bias and a subsequent fine-tunning for knowledge transfer. The past decade of research has clearly demonstrated that understanding the role of texture and shape bias in models like Vision Transformers (ViT) and ResNet, along with the impact of fine-tuning and augmentation techniques, is crucial to enhancing model accuracy and human-like error consistency.
An exploration into shape bias reveals that ViT [149], despite exhibiting a higher shape bias, tends to align more closely with human-like error consistency when fine-tuned. In contrast, ResNet maintains a more human-like error distribution. Augmentation techniques or introduction of shape prior could play a pivotal role in reshaping model representations, impacting their similarity to human classification behavior [151].
While Transformer networks, especially ViT, showcase superior accuracy, they also exhibit a higher shape bias, contrary to most ConvNets which predominantly show texture bias [156, 148, 165, 166]. The fine-tuning of models with augmented data further influences their error consistency, with ViT trending towards a more human-like pattern [166]. The anticipated result could stem from the diminished influence of inductive bias. However, a drawback arises with ViTs [167] as they necessitate a significant volume of data for training to achieve satisfactory quality, contrasting sharply with the considerably lower data requirements in humans.
5.3.5 Humans as Multi-Modal Foundation Models:
There are various reasons to consider humans as complex, natural multi-modal foundation models (FMs), particularly when comparing the biological processes involved in human development with the training technique of any multi-modal agent. Human development dictates a specific development order of multiple sensory modalities such as tactile, sound, vision, and vestibular. At early ages, we learn to combine these inputs to form a coherent understanding of the world, just like how multi-modal foundation models (FMs) integrate data from various sources like images, text, and audio. The trend in multi-modal FM training is that different types of data (text, images, audio) are aligned into a shared representation by treating two modalities as the core pair and align the rest of the modalities through additional trainings/fine-tunnings [168]. However, the order in which the sensory inputs are presented to the model training process does not necessarily align with how the brain integrates multiple sensory inputs into a unified understanding of an object or situation. Though humans can definitely serve as unique inspiration for training multi-modal FMs, as biological processes provide insights into optimizing the selection of modality pairs (anchoring, etc.) and determining the sequence of adding modalities during the training phase.
After the critical period (from birth to around the age of two or three), humans begin developing higher levels of predictive abilities and cognitive functions based on early childhood basic learning and gradually fine-tune their knowledge through experience. This is also true with foundation models that undergo pretraining and fine-tuning gradually upon the improved fidelity and availability of a new modality. In humans, early development builds a broad base of general knowledge, perhaps a common sense of causal events or acquire language through general exposure to words and later experiences help refine skills (predictions, grammar, etc.) and adapt to new environments or tasks, much like how models can be fine-tuned for specific application-required tasks after initial training. In fact, this basic learning phase is the fundamental reason for generalizing knowledge learned from one domain to another. We begin to witness similar emergent behaviors in recent development of FMs [169, 170, 75, 122].
The architectural convergence between biological systems and artificial models is also evident between humans and SoTA multi-modal foundation models. In the human brain, different sensory inputs—such as vision, auditory, and haptic—are processed by distinct neural populations or specialized brain regions (like the occipital lobe specialized for vision and the auditory cortex specialized for sound) [63, 113, 91]. Additionally, even within a single demanding sensory modality, such as vision, separate brain regions handle different aspects like spatial frequency, shape, color, and motion. This separation of sensory processing into distinct pathways mirrors the multi-tower architectures of modern foundation models, where different towers (or specialized subnetworks) handle different types of input data, such as text, images, or audio. For instance, in multi-tower models like CLIP (Contrastive Language–Image Pretraining) [171], one tower processes image data while another handles text, and the two are aligned in a shared latent space. Similarly, transformers are increasingly designed to handle multiple modalities with dedicated structures for each, optimizing the way they learn across diverse data types. Just as the brain integrates separate streams of sensory input to form a unified perception of the environment, multi-modal FMs combine these towers to perform complex cross-modal tasks, such as generating descriptions from images [172] or synthesizing speech from text [173]. This convergence of architecture between biological and computational systems points to a future where AI systems might have the capacity to replicate even more human-like processing efficiencies/capabilities to build general intelligence for general-purpose tasks.
6 Final Thoughts and Discussions
In order to further move forward with our discussions about the strategic placements of humans in the development an AI pipeline, we primarily need to analyze it from where and why all the hype for current AI is coming. To our limited observation, there must be at least two important genres of reality where AI seems to compete with human performance and excel: (1) First, the specifically controlled tasks and/or applications like the ones we see in DeepMind’s game of AlphaGo [174]. The type of AI that accomplishes this specific task is usually referred as “reinforcement learning” and cover only the closed-ended domains of realisms, mostly referring to and exploits a coverage of surface correlations as well as rule-based system of equations, which can be precisely defined (at least to a reasonable degree). This can also be compared to earlier generations (referred as the first generation in some literature [95]) of AI in which symbolic elements and objects are used to define realism and to solve the intelligence problem within that realism. (2) Secondly, lower dimensional realisms or symbolic formal systems which usually manifests itself in the form of human constructs such as natural languages or cave drawings, diverse expressive forms of thoughts, feelings and emotions. We can think of these cognitive models of perception as projections of the external world realism or symbolic formalism which is a multi-dimensional chaotic and complex process that it might become impossible to model. While in some literature, applications are categorized as ones coming from closed-ended domains, such as rule-based games, object classification, and speech recognition, the means and ways our brains develop intelligence towards such a complex realism is popularly called common-sense and can be categorized as open-ended domain.
One observation (which can be formalized into a hypothesis) is that as we go to lower dimensional projections of our realism, objects living within that realism (e.g. words, numbers, concepts, or any meaningful closed form) seem to show the more correlative structure or discernable sensible patterns which naturally become easier to conceptualize and reduce down to the concept of ”correlations”. This has the danger of leading to misunderstandings which might manifest itself in the form of spurious correlations [175]. In other words, to simplify rather complex phenomenon and its representations in the brain, we concisely interfere and make an attempt to reduce the dimensionality to increase the surface correlations. This ultimately lends itself to the absence of common sense, counterfactual reasoning and comprehensive perception –like the 3D or higher dimensional understanding of realism and interactions with it, intertwined with causal reasoning and development of internal models around the external phenomena.
Despite this observation, outstanding performance of the computational frameworks led many research groups to heavily rely on the idea that incorporating human expertise into machine learning systems could be undesirable for introducing bias and, to some extent, dishonesty [176]. In fact, they proposed that a sizable portion of the machine learning community today would be greatly aided if common sense could be inferred from large-scale pleathora of research with the least amount of prior information that could be inherited from a human model. This trend might be the reason for the emergence of self-supervised learning and more broadly self-motivated or self-improving ML [177]. This new trend, in a way I suppose, represents the sheer desire of many AI leaders to search for alternative intelligences that could have better shortcuts compared to evolutionary trajectory of human intelligence development for convergence to an equilibrium: reaching a potentially super-human intelligence.
We could also formulate a hypothesis that bio-inspired AI has a better potential to be more ethical, fair and less judgmental. This line of thinking suggests that AI models derived from biological systems might be better equipped to embody ethical behavior, fairness, and reduced bias compared to other AI forms. When making such a claim, we assume that the architectural and functional properties of a complex system have a direct impact on all the useful emergent behaviors it generates. Drawing inspiration from biological systems, these models may align more closely with ethical principles, exhibiting less judgmental behavior. Several reasons support this stance. The foremost is the comprehensibility and interpretability of such models: enhanced interpretability aids in integrating fairness, ethics, and human values. Biologically plausible or bio-inspired models, grounded in well-explored biological systems, benefit from extensive studies that unravel their functional anatomy. This inherent comprehensibility simplifies the interpretation of their roles in various contexts, thereby facilitating the integration of ethical considerations.
Additionally, human values stem from emergent behaviors within the human brain’s peculiar anatomy and the unique structural characteristics composing the neural system. Therefore, it shouldn’t be surprising to anticipate similar behaviors emerging from bio-inspired systems. This idea holds the view that the behavioural similarity could be coupled with the endogenous structural similarities. While it’s plausible for other forms of intelligence to exhibit these human traits, the likelihood of such emergence might be comparatively lower. In other words, it seems reasonable to hypothize that similar structural properties might lead to similar functional and operational characteristics under certain conditions. The genuine research work is needed to precisely define such conditions and demonstrate the inherent coupling between the hardware (the architecture) and the software (algorithm and/or learnable applications).
Thirdly, bio-inspired systems inherently possess a greater level of inductive bias. This construction, characterized by significant inductive biases, necessitates fewer data samples for learning. While this might appear to limit the overall learning process, human values like fairness or ethical behavior can be guided by a limited number of examples throughout a lifetime. In other words, this might suggest that the innateness could be driven largely by the structural properties. In essence, incorporating these values can be engineered by adjusting the inductive bias and implementing specific structural alterations to the architectural representations. Modifying these aspects in a bio-inspired model should prove simpler compared to an unguided model that approaches the problem from an entirely different intelligence standpoint.
7 Conclusion
Reflecting on the strategic placement of humans within AI development, it becomes critical to examine the origins of the current AI hype. A closer look, as presented in this paper, reveals two distinct domains where algorithms has not only rivaled but, in some cases, surpassed human performance. The first domain is characterized by controlled tasks, such as those mastered by DeepMind’s AlphaGo, where reinforcement learning has thrived within closed-ended systems, exploiting well-defined surface correlations and rule-based frameworks. This echoes the earlier, symbolic AI approaches that tackled intelligence within similarly constrained realisms. The second domain is related to lower-dimensional realities, where human constructs like natural language or symbolic formal systems—often perceived as projections of the external 3D world—simplify complex phenomena into discernible patterns. However, as we transition to higher-dimensional, open-ended domains like common-sense reasoning, the complexity of modeling these realities becomes apparent, highlighting the limitations of current AI methodologies, most starkly the foundation models.
Despite the remarkable advances in computational frameworks mainly inspired by human cognition/perception, there has been a recent growing tendency to minimize human involvement in AI development, driven by growing concerns about introduced bias. This trend has led to the rise of self-supervised learning approaches, to achieve super-human intelligence that could potentially outpace the evolutionary trajectory of human cognition. Yet, amid this shift, bio-inspired AI emerges as a compelling alternative. By drawing on the structural and functional insights from biological systems, bio-inspired AI offers a pathway to more informed, ethical, fair, and interpretable models. These systems, grounded in well-explored biological frameworks, are shown to align with human values inherently and may require fewer data samples to learn ethical behaviors due to their significant and more aligning (with the 3D world model) inductive biases.
The hypothesis that similar structural properties could lead to analogous functional outcomes underscores the potential of bio-inspired AI. The coupling between architectural design and algorithmic functionality in these models suggests a closer alignment with human cognitive processes, offering a more natural integration of ethical, aesthetic, and moral considerations. As the debate continues over the future direction of AI technology, bio-inspired approaches present a promising avenue for bridging the gap between computational efficiency and the human-specific attribute imperatives that must guide the evolution of the efforts for artificial general intelligence.
8 Acknowledgement
The ideas presented in this document stem from a reassessment of my interpretations of the human role in the vast domain of intelligence, informed heavily by local talks at MIT (such as those at the Simons Center for the Social Brain, the Center for Brains, Minds & Machines, etc.), interactions with the faculty at MIT as well as in neighborhood institutions and researchers at vision conferences (VSS, etc.), technical literature readings. In particular, collaborative discussions initiated with experts at the Sinha Lab. (in particular Prof. Pawan Sinha) as well as MIT meseum (in particular Seth Riskin) have enormously enriched my thoughts and understanding of deep and provocative matters of human vision and computational neuroscience. I extend my deepest gratitude and appreciation to all those involved in these enlightening exchanges.